Required Information [The following information apples to the questions displayed below. Washington Countys Board of Representatives Is considering the construction of a longer runway at the county alrport. Currently, the alrport can handle only private alrcraft and small commuter jets. A new, long runway would enable the airport to handle the midsize Jets used on many domestic flights. Data pertinent to the boards decision appear below Cost of acquiring additional land for runway Cost of runway construction Cost of extending perimeter fence Cost of runway lights Annual cost of maintaining new runway Annual incremental revenue from landing fees $85,5ee 29e,e00 15,437 47,800 23,500 62,500 In addition to the preceding data, two other facts are relevant to the decision. First, a longer runway will require a new snowplow, which will cost $190,000. The old snowplow could be sold now for $19,000. The new, larger plow will cost $17,500 more In annual operating costs. Second, the County Board of Representatives belleves that the proposed long runway, and the maJor Jet service It will bring to the county, will Increase economic activity In the community. The board projects that the Increased economic activity will result in $114,000 per year in additional tax revenue for the county In analyzing the runway proposal, the board has decided to use a 10-year time horizon. The countys hurdle rate for capital projects iS 19 percent. Use Appendix A for your reference. (Use appropriate factor(s) from the tables provided.) In analyzing the runway proposal, the board has decided to use a 10-year time horizon. The countys hurdle rate for capltal projects is 19 percent. The County Board of Representatives belleves that if the county conducts a promotional effort costing $29,000 per year, the proposed long runway will result In substantially greater economic development than was projected originally. However, the board is uncertaln about the actual Increase In county tax revenue that will result. Requirec Suppose the board builds the long runway and conducts the promotional campaign. What would the Increase In the countys annual tax revenue need to be In order for the proposed runways Internal rate of return to equal the countys hurdle rate of 19 percent? (Round Intermediate and final answer to the nearest doller amount.) equired increase in tax revenue
Answer #1
1) Initial investment:
Cost of acquiring additional land for runaway $            85,500
Cost of runway construction $         2,90,000
Cost of extending perimeter fence $            15,437
Cost of runway lights $            47,000
Incremental cost of new snowplow = 190000 -19000 = $         1,71,000
Total incremental initial investment $         6,08,937
2) Annual net revenues (excluding additional tax revenue):
Annual incremental revenue from landing fees $            62,500
Additional tax revenue if, new snowplow is bought (resulting from increase in economic activity) $         1,14,000
Annual cost of maintaining new runway (23500+17500) $           -41,000
Annual promotional cost $           -29,000
Annual net revenues (excluding additional tax revenue) $         1,06,500
3) For an IRR of 19%, the NPV should be 0 when discounted
with 19%.
Hence, at IRR = 19%,
608937 = 106500*PVIFA(19,10)+Required increase in tax revenue*PVIFA(19,10) =
Required increase in tax revenue = (608937-106500*4.33893)/4.33893 = $            33,843